what is the vaule of y in the equation x=3y+5
Answer by mananth(16949)

what is the vaule of y in the equation x=3y+5

x=3y+5
The solution for this is to keep y on one side of the equation & remove all terms on the side of y
Add -5 on both sides of the equation to eliminate 5
x-5 = 3y+5 -5

x-5 = 3y
Divide by 3 on both sides of the equation to eliminate 3
x-5 /3 = 3y/3
x-5 / 3 = y